# Are Humans as Brittle as Large Language Models?

### v1

@@ 5 Analysis
For both human annotators and LLMs, we quantify the distributional differences between predictions elicited by each prompt variant and those elicited by the base prompt. A full list of evaluated prompts is provided in Appendix A.3. For all the tasks, we employ the Jensen-Shannon divergence (Lin, 1991; Dagan et al., 1997) as a measure of distributional dissimilarity.
@@ 5.1 RQ1: How do distributional shifts of LLM outputs reveal prompt brittleness?
Figure 3 illustrates the distributional distances between various prompt variants and the corresponding base prompt across the four evaluated tasks. Across all tasks and five LLMs, we find that prompt changes categorized as neutral generally lead to smaller distributional shifts in model outputs compared to those categorized as sensitive. Interestingly, we observe that changes involving alternative label formulations, either semantically or in formats, exert the most important influence on model output distributions. Modifications including adding definitions and requiring justifications or a confidence score, which are grouped under the sensitive category, exhibit a moderate effect.
While these results broadly align with our hypothesis on the categorization of prompt types from the human-grounded aspect, some exceptions are noteworthy. Within the neutral category, changes to the internal order of labels (LabelOrd) induce substantial distributional shifts compared to changes to other prompt aspects. Within the sensitive category, introducing emotional language results in a relatively weaker impact than other modifications. Surprisingly, seemingly subtle changes, such as typos and capitalization of labels, yield measurable distributional shifts, highlighting LLM sensitivity to some surface variations, especially in labels.
@@
To assess whether different LLMs exhibit similar patterns of brittleness across prompt modifications, we compute Spearman’s rank correlation coefficients (Spearman, 1904) between pairs of models, based on the ranking of divergence scores induced by each prompt variant. Heatmaps of pairwise correlations are provided in Appendix A.2. For the emotion classification task, the models show high agreement in their ranking of prompt sensitivities, indicating a shared trend of distributional shifts. Although other tasks display lower rank correlations, the coefficients remain positive, suggesting a generally similar trend to different prompt modifications across models.
@@
Figure 2: Distributional distance between LLM-generated and human annotations across five LLMs. The evaluated tasks are offensiveness rating (a) and emotion classification (b). The x-axis and y-axis denote the prompt variants for human and LLMs respectively. The model name is displayed on the top of each subfigure. Evaluated prompts can be found in Table LABEL:tab:prompt-off for (a) and Table LABEL:tab:prompt-emo for (b).
@@ 5.2 RQ2: Are humans susceptible to prompt brittleness in ways comparable to LLMs?
To investigate whether prompt brittleness also affects human annotators, we turn to our human annotation study, and compare the response distribution divergences across both LLMs and our human annotators These results are presented in Figure 1.
Overall, we find that human divergences are higher overall, but seem to be less dramatically affected by the specifics of instruction variations than LLMs. Across both tasks, human have an average divergence of 0.33, compared to an average LLM divergence of 0.22, but LLMs show a much larger standard deviation of 0.16, compared to 0.04 for humans. This is largely consistent with a state of affairs where human empirical distributions show greater instability, likely due to their smaller sample size, while showing less instruction-conditioned instability (i.e. brittleness) than LLMs.
However, while humans seem to be less brittle than LLMs, we still observe a clear brittleness effect in human annotations. In order to quantify human sensitivity to instruction variations, we employ aggregated $\chi^{2}$-tests to compare each prompt variation to the base prompt, pooled across all instances which we assume to be independent. Table 4 summaries our findings: at a $p<0.05$ level, we find statistically significant differences in human response distributions with LabelOrd-rev, AltLab-gran-2, AltLab-int-2, and Conf for offensiveness rating, and with Typo-label-3, Emo-care, AltLab-keep-2 for emotion classification. Interestingly, the prompt variations for which we find significant differences largely correspond to those which showed the largest distributional differences for LLM responses, with the sole exception of Emo-care, which affected human responses much more than it did LLMs.
@@ 5.3 RQ3: Which types of prompt variants yield similar responses between humans and LLMs?
Our analysis reveals that certain prompt modifications significantly increase the divergence in the distribution between LLM outputs and human annotations. In the offensiveness rating task, for example, the LabelOrd-rev prompt variant leads to notable deviations in outputs generated by LLaMA-3.1-8B, while the Conf variant causes pronounced discrepancies for both Mixtral-8x7B and Mistral-7B. Similarly, in the emotion classification task, the AltLab-keep and Typo-label variants tend to introduce higher divergence from human labels across most evaluated LLMs.
Conversely, we observe stronger alignment between LLM and human output distributions when both are presented with identical prompts, as opposed to mismatched prompt variants. This effect is particularly evident for the AltLab-gran variant in the offensiveness rating task, especially with LLaMA-3.1-8B, LLaMA-3.3-70B, and Mistral-7B. Additionally, in the emotion classification task, the Typo-label variant yields relatively consistent alignment across all LLMs except for the Falcon3-7B model, which exhibits notable divergence.
@@
Offensiveness rating Syns-co-1 $(562.9)$
LabelOrd-rev 607.2607.2*
@@
AltLab-int-2 621.4621.4**
Conf 612.2612.2*
Emotion classification Syns-co-1 $(425.4)$
@@
AltLab-keep-2 766.7766.7***
Conf $(390.4)$
Def-task $(372.9)$
Table 4: Aggregate $\chi^{2}$-test results for human responses. Results reported are aggregated accross all instances, responses for which are assumed to be independent. *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$, (): $p\geq 0.05$
@@ 6 Conclusions and Future Work
Not all our findings are consistent across models. While this variability may reflect differences in LLM training or decoding strategies, we offer practical insights for prompt engineering and emphasize the value of prompt-aware annotation.
@@
While we explore a wide range of prompt variations, only a limited subset can be evaluated through human annotation due to practical constraints. In particular, the sample size of instances in our human study is not sufficiently large to provide a comprehensive understanding of prompt brittleness across tasks. Expanding the number of tasks and instances would improve the robustness of our results. Among the various types of prompt perturbations investigated, some categories remain underexplored. For instance, in the case of the LabelOrd modification, we do not test all possible permutations of label orderings, which could limit the scope of our conclusions.
Additionally, the computation of distance scores is sensitive to the size of the annotation distribution. Our comparisons between LLM-generated and human annotations could benefit from a more refined calibration strategy to address the discrepancy in the number of annotations per instance between the two sources.
@@
ChatGPT (OpenAI, 2025) was used as a tool to improve code generation for figures and tables, as well as to assist with grammar and vocabulary in the text of some sections of this paper.
@@ Acknowledgments
This work is funded by the project INPROMPT (Interactive Prompt Optimization with the Human in the Loop for Natural Language Understanding Model Development and Intervention, funded by the German Research Foundation, KL 2869/13-1, project number 521755488).
@@ Appendix A Appendix
A.1 Heatmaps of Distance Scores between Prompt Variants
See Figure 3.
@@ A.1 Heatmaps of Distance Scores between Prompt Variants
Figure 3: Heatmaps showing distance scores between the distributions of LLM predictions with prompt variants and with the base prompt for four evaluated tasks in Table 3. Distance scores are calculated using Jensen-Shannon divergence (Dagan et al., 1997). The x-axis represents different types of prompt modifications defined in Section 3, with numeric indices representing multiple instances of the same type-variant pair. The y-axis lists the five LLMs evaluated in the study. Black dashed lines divide each heatmap into three parts: prompt modifications belonging to the neutral class (left), those from the sensitive class (middle), and the average distance scores for neutral and sensitive class modifications respectively (right). All prompts evaluated are provided in Appendix A.3.
A.2 Heatmaps of SpearmanR Statistics Across LLMs
See Figure 4.
(a) Offensiveness rating.
(b) Politeness rating.
(c) Emotion classification.
(d) Irony detection.
Figure 4: Heatmaps showing Spearman’s rank correlation coefficients of distance scores for prompt perturbations across five LLMs. The four evaluated tasks are (a) offensiveness rating, (b) politeness rating, (c) emotion classification, and (d) irony detection. We present the results for prompt perturbations in the neutral and sensitive categories introduced in Section 3. Labels of x-axis and y-axis both refer to the names of LLMs. The coefficient values have a range from -1 to 1, where 1 means perfect monotonic increasing correlation (ranks agree exactly), -1 means monotonic decreasing correlation (ranks are opposites), 0 means no monotonic relationship.
A.3 Examples for Prompt Modifications
@@
A.4 Instructions of Survey for Human Study
@@
Figure 5: A screenshot of the instruction page for offensiveness rating survey. The contact information is blinded.
Figure 6: A screenshot of the instruction page for the emotion classification survey. The contact information is blinded.

### v2

@@ 5 Analysis
For both human annotators and LLMs, we quantify the distributional differences between predictions elicited by each prompt variant and those elicited by the base prompt. A full list of evaluated prompts is provided in Appendix A.2. For all the tasks, we employ the Jensen–Shannon divergence (61115; dagan-etal-1997-similarity) as a measure of distributional dissimilarity to display the human or LLM brittleness to prompt changes.
@@ 5.1 RQ1: How do distributional shifts of LLM outputs reveal prompt brittleness?
Figure 1 illustrates the distributional distances between various prompt variants and the corresponding base prompt across the four evaluated tasks. Across all tasks and five LLMs, we find that prompt changes categorized as neutral generally lead to smaller distributional shifts in model outputs compared to those categorized as sensitive. Interestingly, we observe that changes involving alternative label formulations, either semantically or in substitute mapping formats, exert the most important influence on model output distributions. Modifications including adding definitions and requiring justifications or a confidence score, which belong to the sensitive category, exhibit a moderate effect.
While these results broadly align with our hypothesis on the categorization of prompt types from the human-grounded aspect, some exceptions are noteworthy. Within the neutral category, changes to the internal order of labels (LabelOrd) induce substantial distributional shifts compared to changes to other prompt aspects. Within the sensitive category, introducing emotional language results in a relatively weaker impact than other modifications. Seemingly subtle changes, such as typos and capitalization of labels, yield measurable distributional shifts, highlighting LLM sensitivity to some surface variations, especially in labels.
@@
To assess whether different LLMs exhibit similar patterns of brittleness across prompt modifications, we compute Spearman’s rank correlation coefficients (spearman04) between pairs of models, based on the ranking of divergence scores induced by each prompt variant. Heatmaps of pairwise correlations are provided in Appendix A.1. For the emotion classification task, the models show high agreement in their ranking of prompt sensitivities, indicating a shared trend of distributional shifts. Although other tasks display lower rank correlations, the coefficients remain positive, suggesting a generally similar trend to different prompt modifications across models.
@@
Figure 3: Distributional distance between LLM-generated and human annotations across five LLMs. The evaluated tasks are (a) offensiveness rating and (b) emotion classification. The x-axis and y-axis denote the prompt variants for human and LLMs respectively. The model name is displayed on the top of each subfigure. Evaluated prompts can be found in Table LABEL:tab:prompt-off for (a) and Table LABEL:tab:prompt-emo for (b).
@@ 5.2 RQ2: Are humans susceptible to prompt brittleness in ways comparable to LLMs?
To investigate whether prompt brittleness also affects human annotators, we turn to our human annotation study, and compare the response distribution divergences across both LLMs and our human annotators. These results are presented in Figure 2.
Overall, we find that human divergences are higher for the same prompt, but seem to be less dramatically affected by the specifics of instruction variations than LLMs. Across both tasks, humans have an average divergence of 0.32, compared to an average LLM divergence of 0.22, but LLMs show a much larger standard deviation of 0.14, compared to 0.03 for humans. This is largely consistent with a state of affairs where human empirical distributions show greater instability, likely due to their smaller sample size, while showing less instruction-conditioned instability (i.e. brittleness) than LLMs.
However, while humans seem to be less brittle than LLMs, we still observe a clear brittleness effect in human annotations. In order to quantify human sensitivity to instruction variations, we employ aggregated $\chi^{2}$-tests to compare each prompt variation to the base prompt, pooled across all instances which we assume to be independent. Table 4 summaries our findings: at a $p<0.05$ level, we find statistically significant differences in human response distributions with LabelOrd-rev, AltLab-gran-2, AltLab-int-2, Def-both, Conf, and Def-both for the offensiveness rating task, and with Typo-label-3, Emo-care, AltLab-keep-2 for the emotion classification task.
Interestingly, the prompt variations for which we find significant differences largely correspond to those which showed the largest distributional differences for LLM responses, with the sole exception of Emo-care, which affected human responses much more than it did LLMs. Moreover, although human annotation results remain consistent across both tasks, certain prompt variants (e.g., Cap-label, Def-both, and Conf) elicit markedly greater sensitivity in LLM performance on the offensiveness rating task than on the emotion classification task.
@@ 5.3 RQ3: Which types of prompt variants yield similar responses between humans and LLMs?
Our analysis reveals that certain prompt modifications significantly increase the divergence in the distribution between LLM outputs and human annotations. In the offensiveness rating task, for example, the LabelOrd-rev prompt variant leads to notable deviations in outputs generated by Mixtral-8x7B and Mistral-7B, while the AltLab-gran-2 variant causes pronounced discrepancies for the Llama models. Similarly, in the emotion classification task, the AltLab-keep-2 and Typo-label-3 variants tend to introduce higher divergence from human labels across most evaluated LLMs. Surprisingly, the Exp variant demonstrates that requesting an explanation during annotation enhances the alignment between LLMs and humans.
Conversely, we observe stronger alignment between LLM and human output distributions when both are presented with identical prompts, as opposed to mismatched prompt variants. This effect is particularly evident for the AltLab-gran-2 and Def-both variants in the offensiveness rating task, especially with LLaMA-3.1-8B, LLaMA-3.3-70B, and Mistral-7B. Additionally, in the emotion classification task, the Typo-label and Exp variants yield relatively consistent alignment across almost all evaluated LLMs, which exhibits less divergence.
@@
Offensiveness rating LabelOrd-rev 607.2607.2*
Syns-co-1 $(562.9)$
Typo-label-1 $(511.3)$
Cap-label $(529.4)$
AltLab-int-2 621.4621.4**
@@
Emotion classification LabelOrd-shuff1 $(360.8)$
Syns-co-1 $(425.4)$
@@
Table 4: Aggregate $\chi^{2}$-test results for human responses. In each case, we compare the human response distribution for a prompt variation against the response distribution for the base prompt, in order to test which prompt variations affected human responses in a statistically significant manner. Results reported are aggregated accross all instances, responses for which are assumed to be independent. *: $p<0.05$, **: $p<0.01$, ***: $p<0.001$, (): $p\geq 0.05$
@@ 6 Conclusions and Future Work
Not all our findings are consistent across models. While this variability may reflect differences in LLM training or model sensitivity to the stochastic decoding strategy, we offer practical insights for prompt engineering and emphasize the value of prompt-aware annotation.
@@
While we explore all proposed categories of prompt variations, only a subset of samples within each category can be evaluated through human annotation due to practical constraints. Expanding the number of tasks and instances would further improve the robustness of our results. Among the various types of prompt perturbations investigated, some variants remain underexplored. For instance, in the case of the LabelOrd modification, we do not test all possible permutations of label orderings, which could limit the scope of our conclusions.
Additionally, the computation of distance scores using Jensen–Shannon divergence is sensitive to the size of the annotation distribution. In this study, the human annotations comprise only 10 samples, whereas the LLM-generated annotations comprise 100 samples. The relatively small human annotation set can result in higher variance in the distribution, which may increase the measured divergence compared to the larger LLM annotation set. Our comparisons between LLM-generated and human annotations could benefit from a more refined calibration strategy to address the discrepancy in the number of annotations per instance between the two sources.
@@
ChatGPT (chatgpt) was used as a tool to improve code generation for figures and tables, as well as to assist with grammar and vocabulary in the text of some sections of this paper.
@@ Acknowledgments
This work is funded by the project INPROMPT (Interactive Prompt Optimization with the Human in the Loop for Natural Language Understanding Model Development and Intervention, funded by the German Research Foundation, KL 2869/13–1, project number 521755488).
@@ Appendix A Appendix
A.1 Heatmaps of SpearmanR Statistics Across LLMs
We compute Spearman’s rank correlation coefficients between pairs of models for the four evaluated tasks, based on the ranking of Jensen–Shannon divergence scores induced by each prompt variant. See Figure 4.
@@ A.1 Heatmaps of SpearmanR Statistics Across LLMs
Figure 4: Heatmaps showing Spearman’s rank correlation coefficients of distance scores for prompt perturbations across five LLMs. The four evaluated tasks are (a) offensiveness rating, (b) politeness rating, (c) emotion classification, and (d) irony detection. We present the results for prompt perturbations in the neutral and sensitive categories introduced in Section 3. Labels of x-axis and y-axis both refer to the names of LLMs. The coefficient values have a range from 1-1 to 11, where 11 means perfect monotonic increasing correlation (ranks agree exactly), -11 means monotonic decreasing correlation (ranks are opposites), 0 means no monotonic relationship.
A.2 Examples for Prompt Modifications
@@
A.3 Instructions of Survey for Human Study
@@
Figure 5: A screenshot of the instruction page for offensiveness rating survey.
Figure 6: A screenshot of the instruction page for the emotion classification survey.